To test Ho : ?= 20 versus H1 : ?< 20, a simple random sample of size n = 17 is obtained from a population that is known to be normally distributed Answer parts (a)-(d) E Click here to view the t-Distribution Area in Right Tail (a) If x 18.3 and s 3.8, compute the test statistic. t-(Round to two decimal places as needed.)
Answer #1

The test statistic can be written as

Vn(T- 20 which under H0 follows a t distribution with n-1 df.

Here

The value of the test statistic t_{obs} = \frac{\sqrt{17}(18.3-20)}{3.8} = -1.84455 (ans)